OpenStudy (anonymous):

Write the function in slope-intercept form. -5x + y = 7

OpenStudy (anonymous):

y=mx+b form, use some algebra

OpenStudy (anonymous):

-5x+y = 7 +5x +5x y = 5x+7 done

OpenStudy (anonymous):

thank you @iambatman i am so horrible with math can you help me with this one? Identify the x-intercept and the y-intercept for the line. 3x + 2y = -6

OpenStudy (anonymous):

I'm sure you're not bad, just requires some effort, to find the x - intercepts set y = 0, and to find y - intercept set x = 0, and evaluate.